I can set up the problem but I can’t seem to get the right answer. Please show me what I am doing wrong?
Problem: The Alvin, submersiblue research vessel, can descend into the ocean to a depth of approximately 4500 m in just over 5hr. Determine its average rate of submersion in ft per min and km per second.
What I did: 4500 m/5hr x 1hr/60min x 5280ft/1m= 79200.000 ft/min

4500 m/5ft x 1hr/60min x 1min/60sec x 1000km/1m= 250.000 km/sec
The correct answers: 50ft/min and 3 x 10 to the -4 power km/sec

Found 3 solutions by rothauserc, ikleyn, greenestamps:
Answer by rothauserc(4718)

rate of submersion in ft / m
:
5 hr * 60 = 300 min
:
1 m = 3.28 ft
:
4500 * 3.28 = 14760 ft
:
***********************************************
14760 / 300 = 49.2 is approximately 50 ft / min
***********************************************
:
rate of submersion in km / s
:
4500 m = 4.5 km
:
5 hr = 5 * 60 * 60 = 18000 seconds
:
******************************************************
4.5 / 18000 = 0.00025 is approximately 3E10^-4 km / s
******************************************************
:

Answer by ikleyn(52800)

.

            See how the other tutor makes conversion in his post and learn from there.

            The way on how you tried to do it in your post will only lead you to numerous errors.

            When you make conversion,  always think on how to do it by a most simple way.


===================

Added after reading the post by @greenestamps.

    I only can repeat it AGAIN from my post above


        The way on how you tried to do it in your post is formally correct, by in practice will only lead you to numerous errors.


         And what is written in the post by @greenestamps, only confirms my statement.

For many years working at the engineering position in Civil Engineering in Unated States, I performed units conversions many times per day
7 days per week (among other my duties) and KNOW what I am saying.


Answer by greenestamps(13200)



The method you are using -- multiplying the given measurement by unit fractions to convert the units without changing the measurement -- is a perfectly good way to work problems like this; in some similar problems where the units are unfamiliar, it is by far the easiest method to use.

But you have to use the correct unit fractions....
   4500m    1 hr    3.28 ft   49.2 ft
   ----- * ------ * ------- = -------
    5 hr   60 min     1 m      1 min

(1 meter = 3.28 feet, not 5280 feet)
   4500m    1 hr     1 min    1 km    2.5*10^(-4) km
   ----- * ------ * ------ * ------ = --------------
   5 hr    60 min   60 sec   1000 m        1 sec

(1 km is 1000 meters; not the other way around)

----------------------------------------------------------------

Added in response to the post from tutor @ikleyn, in response to this response from me....

Dear tutor @ikleyn....

Why do you have to think the way you work problems is right?!

What about the method of unit fractions leads to errors? The only way to make an error with that method is to use it improperly. The same can be said of any other method for solving this kind of problem, or any kind of problem.

What in my response "confirms" that the method of unit fractions leads to errors? My results are more accurate than the "correct" answers given in the posted problem.

I too have a huge amount of experience in performing unit conversions; and I also have a huge amount of experience seeing students performing unit conversions. In my working with students, the method of unit conversions is far more successful than any other.

Please keep to your usually excellent responses to questions, and stop criticizing other responses from other tutors just because they don't agree with the way you do things.